OpenStudy (anonymous):

(sin x / 1 - cos x ) + ( sin x / 1+ cos x ) = 2 csc x

OpenStudy (anonymous):

Verify each trigonometric equation by substituting identities to match the right hand side of the equation to the left hand side of the equation.

OpenStudy (ankit042):

six(1+cosx+1-cosx)/(1-cos^2x) 2sinx(sin^2x) 2/sinx

OpenStudy (anonymous):

is that the whole answer ? @ankit042

OpenStudy (anonymous):

\[\frac{ \sin(x) }{ 1-\cos(x) }+\frac{ \sin(x) }{ 1+\cos(x) }=\frac{ 2\sin(x) }{ 1-\cos^2(x)}=\frac{ 2\sin(x) }{ \sin^2(x) }=\frac{ 2 }{ \sin(x) }=2cosec(x)\]

OpenStudy (ankit042):

yes 2/sinx = 2cosecx